OpenStudy (anonymous):

MEDAL A company sells boxes of bracelets (b) for $30 and boxes of necklaces (n) for $50. They make batches of bracelets that fill 5 boxes and batches of necklaces that fill 7 boxes. The company only has 48 boxes. Which system of equations helps the company plan to make $320? b = 48 - n 5b + 7n = 48 b = 48 - n 30b + 50n = 320 5b = 48 - 7n 30b + 50n = 320 5b = 48 - 7n 30b - 50n = 320

OpenStudy (anonymous):

b = 48 - n 5b + 7n = 48 b = 48 - n 30b + 50n = 320 5b = 48 - 7n 30b + 50n = 320** this one! <====== ;D 5b = 48 - 7n 30b - 50n = 320